Grade 12Algebra

Use the sign of >,
– 231 + 79 + 51 [ ] -399 + 159 + 81

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

Let us take Left Hand Side (LHS) = - 231 + 79 + 51
= - 231 + 130
= -101
Now,
Right Hand Side (RHS) = - 399 + 159 + 81
= - 399 + 240
= - 159
By comparing LHS and RHS
LHS > RHS
-101 > -159
∴–231 + 79 + 51 [>] -399 + 159 + 81

Thanks